OpenStudy (tinathetypical):

Given the equation: 5y+24+2xi=10x-10yi-15i, find x and y if there are both real numbers

OpenStudy (tinathetypical):

please help!!!

OpenStudy (tkhunny):

Equality of complex numbers requires two pieces. a + bi = c+di iff a = c and b = d Do that.